# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 2.3. The $\varphi_{\lambda}$-transformation and the null transformation
For any fixed $\lambda\in\mathbb{C}$, the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ yield a transformation
@@
We call it the $\varphi_{\lambda}$-transformation of $(y^{1},\dots,y^{\ell})$. In particular, the $\varphi_{0}$ transformation is called the null transformation.
@@
The corresponding null transformation is given by
@@
The Jacobian determinant of the null transformation is given by
$\mathbf{J}_{0}(\hat{\mathbf{x}})=e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+},\,(\beta,\omega)=0}\left(1-\mathrm{e}^{-2\pi i(\beta,\hat{\mathbf{x}})}\right),$ (2.21)
where $\Phi^{+}$ is the set of all positive roots of $R$.
@@
$\mathbf{J}_{\lambda}(\hat{\mathbf{x}})~{}=$ $~{}\mathrm{e}^{-2\pi ic\sum_{j}(\omega_{j},\omega)+\pi i\sum_{\alpha\in\Phi^{+}}\left(c(\alpha,\omega)+(\alpha,\hat{\mathbf{x}})\right)}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right)$
$=$ $~{}e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right).$ (2.22)
@@
$\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha=\sum_{j=1}^{\ell}\omega_{j}.$ (2.23)
Since $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is a polynomial of $\lambda$, its constant term w.r.t. $\lambda$ yields (2.21). The proposition is proved. ∎
@@
The Jacobian determinant of the null transformation is non-vanishing except on an $(\ell-1)$-dimensional divisor. In particular, $y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0}$ are functionally independent.
Theorem 2.6.
We have $\mathscr{A}^{W}=\mathbb{C}[y^{1},\dots,y^{\ell};\lambda]$.
Proof.
For any given $f\in\mathscr{A}^{W}$, we need to show that it can be represented in terms of a polynomial of $y^{1},\dots,y^{\ell}$ and $\lambda$. By using Lemma 2.1, we know that $f$ can be decomposed as the sum of homogeneous elements of $\mathscr{A}^{W}$, so we may assume without lose of generality that $f$ is homogeneous of degree $n$.
From [6] we know that the Fourier polynomial $f|_{\lambda=1}$ of $x^{1},\dots,x^{\ell}$ can be represented in terms of a polynomial of $y^{1}\big{|}_{\lambda=1}$,…,$y^{\ell}\big{|}_{\lambda=1}$ as follows:
@@
so it follows from the homogeneity of $f$ and of $y^{1},\dots,y^{\ell}$ that
@@
here $n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}\in\mathbb{Z}$. Hence we can represent $f$ in the form
@@
This leads to a contradiction with the fact, as shown by Corollary 2.5, that $y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0}$ are functionally independent. Thus $q$ must be a non-negative integer, and the proposition is proved. ∎
Definition 2.2.
We call $\{y^{1},\dots,y^{\ell}\}$ the set of basic generators of $\mathscr{A}^{W}$.
@@ 2.4. Leading terms of the basic generators
which we call the leading term of $y^{j}$. This term also appears in $y^{j}|_{\lambda=0}$, and the following theorem shows that for some indices $j$, $y^{j}|_{\lambda=0}$ only contain their leading terms.
@@
Let us denote
$S=\{r\in\{1,\dots,\ell\}\mid m_{r}=(\omega,\alpha_{r}^{\vee})>0\}.$ (2.29)
@@
$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ (\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (2.31)
@@
$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0,$ (2.32)
@@
$(\sigma^{*}(\omega_{j}),\omega_{j})=(\omega_{j},\omega_{j})-\sum_{r=1}^{\ell}n_{r}(\alpha_{r},\omega_{j})=(\omega_{j},\omega_{j}).$ (2.33)
Since $\sigma^{*}$ is an orthogonal transformation, the equation (2.33) implies that $\sigma^{*}(\omega_{j})=\omega_{j}$. Therefore, from (2.31) it follows that
$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi i(\omega_{j},\hat{\mathbf{x}})}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi ix^{j}}.$ (2.34)
@@
$\sum_{r=1}^{\ell}m_{r}\log\left(y^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.35)
2.5. Proper generators
In order to construct generalized Frobenius manifold structures on the orbit space of $W_{a}(R)$, we introduce in this subsection some other generators of $\mathscr{A}^{W}$.
@@ 2.5. Proper generators
$\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell},\lambda].$ (2.38)
@@
It suffices to show that $y^{j}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for $j=1,\dots,\ell$. Assume that there exist some $y^{j}$ which do not belong to $\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, and let $y^{k}$ be one of those with the lowest degree. It follows from (2.37) that $s^{k}\not\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$. However, since $\theta_{k}$ is the lowest degree among the ones of $y^{j}\not\in C[z^{1},\dots,z^{\ell},\lambda]$, we know that $y^{r}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for each $r$ satisfying $\theta_{r}<\theta_{k}$. Then we know from Lemma 2.9 that $s^{k}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, which leads to a contradiction. The proposition is proved. ∎
@@
$\det\left(\frac{\partial z^{j}}{\partial x^{p}}\right)=\det\left(\frac{\partial y^{j}}{\partial x^{p}}\right)=\mathbf{J}_{\lambda}(\hat{\mathbf{x}}),$ (2.39)
where $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is defined in (2.22), thus $z^{1},\dots,z^{\ell}$ are functionally independent for each $\lambda\in\mathbb{C}$.
@@
For each $1\leq j\leq\ell$ we have $z^{j}=y^{j}+\lambda s^{j}$, where $s^{j}$ is a certain polynomial of $\lambda$ and $y^{r}$ satisfying the condition $\theta_{r}<\theta_{j}$. Then from the relation
@@
For any given set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we will also call the transformation which maps $(x^{1},\dots,x^{\ell})$ to $(z^{1},\dots,z^{\ell})$ the $\varphi_{\lambda}$-transformation, and the corresponding $\varphi_{0}$-transformation the null transformation. Note that $\{z^{1},\dots,z^{\ell}\}$ and $\{y^{1},\dots,y^{\ell}\}$ may lead to distinct $\varphi_{\lambda}$-transformations, but they always share the same null-transformation, i.e.,
In particular, for any element $r$ of the set $S$ defined by (2.29) we have
@@ 3. The construction of generalized Frobenius manifold structures
In this section, we are to construct a generalized Frobenius manifold structure starting from a given irreducible reduced root system $R$ in the $l$-dimensional Euclidean space $V$, a fixed weight $\omega$ of the form (2.1), and a set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$ satisfying a certain additional condition.
@@ 3.1. The orbit space and pencil generators
By definition, $\mathbb{C}[z^{1},\cdots,z^{\ell}]$ is a subalgebra of $\mathscr{A}^{W}$. Denote the spectrum of this subalgebra by $\mathcal{M}$, then $\mathcal{M}\cong\mathbb{C}^{\ell}$. Taking $x^{1},\cdots,x^{\ell}$ as coordinates on the complexified Euclidean space $V\otimes\mathbb{C}$, we have a collection of holomorphic maps
@@
Here the $\varphi_{\lambda}$-transformation is defined in the last section. We call $\mathcal{M}$ the orbit space of the affine Weyl group $W_{a}(R)$ w.r.t. the proper generators $z^{1},\dots,z^{\ell}$. In this section, we will use this collection of maps $\{\varphi_{\lambda}\}_{\lambda\in\mathbb{C}}$ to endow $\mathcal{M}$ with a structure of generalized Frobenius manifold.
[…]
$a(\mathrm{d}x^{i},\mathrm{d}x^{j})=a^{ij}\ \mbox{with}\ (a^{ij})=\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)^{-1},\quad i,j=1,\dots,\ell$ (3.2)
on $V\otimes\mathbb{C}$. Here by a contravariant metric we mean a symmetric non-degenerate bilinear form on the holomorphic cotangent bundle $T^{*}(V\otimes\mathbb{C})$.
@@
Note that we introduce the factor $1/{4\pi^{2}}$ to simplify the expressions of $g_{\lambda}^{ij}$’s.
@@
We have $g^{ij}_{\lambda}\in\mathscr{A}^{W}$.
@@
By definition, we have $g^{ij}_{\lambda}\in\mathscr{A}$, hence it suffices to show that $g^{ij}_{\lambda}$ is invariant w.r.t. the action of $W_{a}(R)$. Since the invariance of $g^{ij}_{\lambda}$ w.r.t. translations is obvious, we consider actions by elements of the Weyl group $W(R)$. Write $z^{j}$ as $z^{j}=z^{j}(x^{1},\dots,x^{\ell};c)$. For any given $\sigma\in W(R)$, since $z^{j}(x^{1},\dots,x^{\ell};c)=z^{j}(\sigma(x^{1},\dots,x^{\ell};c))$, we have
@@
$\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}=a^{pq}.$
Hence $g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=g^{ij}_{\lambda}(\sigma(x^{1},\dots,x^{\ell};c))$. The proposition is proved. ∎
The above proposition shows in particular that $g^{ij}_{\lambda}$ are polynomials of $\lambda$. In order to construct generalized Frobenius manifold structures on the orbit space $\mathcal{M}$, we are to impose an additional condition on the proper generators of $\mathscr{A}^{W}$, so that $g_{\lambda}$ yields a flat pencil of metrics [10] on $\mathcal{M}$.

### v3

@@ 2.3. The $\varphi_{\lambda}$-transformation and the null-transformation
For any fixed $\lambda\in\mathbb{C}$, the basic generators $y^{1},\dots,y^{\ell}$ yield a transformation
@@
We call it the $\varphi_{\lambda}$-transformation of $(y^{1},\dots,y^{\ell})$. In particular, the $\varphi_{0}$ transformation is called the null-transformation.
@@
The corresponding null-transformation is given by
@@
The Jacobian determinant of the null-transformation is given by
[…]
where $R^{+}$ is the set of all positive roots of $R$ and $C$ is a nonzero constant.
@@
[…]
@@
$\frac{1}{2}\sum_{\alpha\in R^{+}}\alpha=\sum_{j=1}^{\ell}\omega_{j}.$ (2.23)
Since $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ is a polynomial of $\lambda$, we have $\mathbf{J}_{0}(\hat{\mathbf{x}})=\mathbf{J}(\hat{\mathbf{x}};0)$, so the constant term of $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ w.r.t. $\lambda$ yields (2.21). The proposition is proved. ∎
@@
The Jacobian determinant of the null-transformation is non-vanishing except on an $(\ell-1)$-dimensional divisor. In particular, $y^{1}\big|_{\lambda=0},\dots,y^{\ell}\big|_{\lambda=0}$ are functionally independent.
We now turn to the proof of Theorem 2.4.
Proof of Theorem 2.4.
For any given $f\in\mathscr{A}^{W}$, we need to show that it can be represented in terms of a polynomial of $y^{1},\dots,y^{\ell}$ and $\lambda$. By using Lemma 2.1, we know that $f$ can be decomposed as the sum of quasi-homogeneous elements of $\mathscr{A}^{W}$, so we may assume without lose of generality that $f$ is quasi-homogeneous of degree $n$.
From [6] we know that the Fourier polynomial $f|_{\lambda=1}$ of $x^{1},\dots,x^{\ell}$ can be represented in terms of a polynomial of $y^{1}\big|_{\lambda=1}$$y^{\ell}\big|_{\lambda=1}$ as follows:
@@
so it follows from the quasi-homogeneity of $f$ and of $y^{1},\dots,y^{\ell}$ that
@@
where each $n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}\in\mathbb{Z}$. Hence we can represent $f$ in the form
@@
This leads to a contradiction with the fact, as shown by Corollary 2.7, that $y^{1}\big|_{\lambda=0},\dots,y^{\ell}\big|_{\lambda=0}$ are functionally independent. Thus $q$ must be a non-negative integer, and the theorem is proved. ∎
@@ 2.4. Leading terms of the basic generators
which we call the leading term of $y^{j}$. Since not involving $\lambda$, this term also appears in $y^{j}|_{\lambda=0}$. The following theorem shows that for some indices $j$, $y^{j}|_{\lambda=0}$ consist merely of their leading terms.
@@
Denote
$S=\{r\in\{1,\dots,\ell\}\colon m_{r}=(\omega,\alpha_{r}^{\vee})>0\},$ (2.29)
@@
$y^{j}\big|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W(R)\\ (\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\hat{\mathbf{x}})}$. (2.31)
@@
$(\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)=0,$ (2.32)
@@
$(\sigma^{-1}(\omega_{j}),\omega_{j})=(\omega_{j},\omega_{j})-\sum_{r=1}^{\ell}n_{r}(\alpha_{r},\omega_{j})=(\omega_{j},\omega_{j}).$ (2.33)
Since $\sigma^{-1}$ is an orthogonal transformation, the equation (2.33) implies that $\sigma^{-1}(\omega_{j})=\omega_{j}$. Therefore, from (2.31) it follows that
$y^{j}\big|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W(R)\\ \omega_{j}=\sigma^{-1}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi i(\omega_{j},\hat{\mathbf{x}})}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W(R)\\ \omega_{j}=\sigma^{-1}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi ix^{j}}.$ (2.34)
@@
[…]
2.5. Proper factors and proper generators
In order to construct generalized Frobenius manifold structures on the orbit space of $W_{a}(R)$, we introduce in this subsection the notion proper generators and proper factors of $\mathscr{A}^{W}$.
Consider the subalgebra $\mathscr{R}=\mathbb{C}[y^{1},\dots,y^{\ell}]$ of $\mathscr{A}^{W}$. It has the following two properties:
  • $\mathscr{R}$ is a graded subring of $\mathscr{A}$. This means that for any $f\in\mathscr{R}$, if we decompose it into quasi-homogeneous components as $f_{1}+\cdots+f_{n}$, then each $f_{k}\in\mathscr{R}$. This property follows from the quasi-homogeneity of $y^{1},\dots,y^{\ell}$.
  • The tensor product decomposition $\mathscr{A}^{W}=\mathscr{R}\otimes\mathbb{C}[\lambda]$ holds. This means that the natural homomorphism
    $h\colon\mathscr{R}\otimes\mathbb{C}[\lambda]\to\mathscr{A}^{W},\quad f\otimes g\mapsto fg,\quad f\in\mathscr{R},\,g\in\mathbb{C}[\lambda]$
    is an isomorphism.
These properties give rise to the following definition.
Definition 2.10.
A subalgebra $\mathscr{P}$ of $\mathscr{A}^{W}$ is called a proper factor, if it is graded and $\mathscr{A}^{W}=\mathscr{P}\otimes\mathbb{C}[\lambda]$.
In what follows, we are to find a set of generators for each proper factor $\mathscr{P}$, called proper generators.
@@ 2.5. Proper factors and proper generators
$\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell};\lambda].$ (2.38)
@@
It suffices to show that $y^{j}\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$ for $j=1,\dots,\ell$. Assume that there exist some $y^{j}$ which do not belong to $\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, and let $y^{k}$ be one of those with the lowest degree. It follows from (2.37) that $s^{k}\not\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$. However, since $\theta_{k}$ is the lowest degree among the ones of $y^{j}$ that do not belong to $C[z^{1},\dots,z^{\ell};\lambda]$, we know that $y^{r}\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$ for each $r$ satisfying $\theta_{r}<\theta_{k}$. Then we know from Lemma 2.12 that $s^{k}\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, which leads to a contradiction. The proposition is proved. ∎
@@
$\det\left(\frac{\partial z^{j}}{\partial x^{p}}\right)=\det\left(\frac{\partial y^{j}}{\partial x^{p}}\right)=\mathbf{J}(\hat{\mathbf{x}};\lambda),$ (2.39)
where $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ is defined by (2.22). Thus $z^{1},\dots,z^{\ell}$ are functionally independent for each $\lambda\in\mathbb{C}$.
@@
For each $1\leq j\leq\ell$ we have $z^{j}=y^{j}+\lambda s^{j}$, where $s^{j}$ is a certain polynomial of $\lambda$ and $y^{r}$ that satisfy the condition $\theta_{r}<\theta_{j}$. Thus from the relations
@@
Suppose $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators, then $\mathscr{P}=\mathbb{C}[z^{1},\dots,z^{\ell}]$ is a proper factor. In fact, it follows from the corollary above that $z^{1},\dots,z^{\ell}$ and $\lambda$ are algebraically independent (as functions of $x^{1},\dots,x^{\ell}$ and $\lambda$), hence $\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell}]\otimes\mathbb{C}[\lambda]$. Since $z^{1},\dots,z^{\ell}$ are quasi-homogeneous, the subalgebra $\mathbb{C}[z^{1},\dots,z^{\ell}]$ is graded. The following theorem shows that all proper factors have such a form. Furthermore, each proper factor is generated by a unique set of proper generators.
Theorem 2.15.
Suppose $\mathscr{P}$ is a proper factor of $\mathscr{A}^{W}$, then there exists a unique set of proper generator $\{z^{1},\dots,z^{\ell}\}$ such that $\mathscr{P}=\mathbb{C}[z^{1},\dots,z^{\ell}]$.
Proof.
By definition, we have $\mathscr{A}^{W}=\mathscr{P}\otimes\mathbb{C}[\lambda]$, hence each $y^{i}\in\mathscr{A}^{W}$ can be represented in the form
$y^{i}=\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(\lambda),\quad p^{i}_{\alpha}\in\mathscr{P},\ T^{i}_{\alpha}(\lambda)\in\mathbb{C}[\lambda].$ (2.41)
In particular, we have
$y^{i}\big|_{\lambda=0}=\Bigl(\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(0)\Bigr)\Bigl|_{\lambda=0},$
where $\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(0)\in\mathscr{P}$. So we define $z^{i}$ to be the $\theta_{i}$-degree component of $\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(0)$, then $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators. Since $\mathscr{P}$ is graded, we have $z^{i}\in\mathscr{P}$, thus $\mathbb{C}[z^{1},\dots,z^{\ell}]\subset\mathscr{P}$. However, it follows from the properties of proper generators that $\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell}]\otimes\mathbb{C}[\lambda]$. Since $\mathbb{C}[\lambda]$ is faithfully flat, we have
$\mathbb{C}[z^{1},\dots,z^{\ell}]=\mathscr{P}.$
Now, let us prove the uniqueness of the set of proper generators. Suppose there is another set of proper generators $\{\tilde{z}^{1},\dots,\tilde{z}^{\ell}\}$ such that $\mathscr{P}=\mathbb{C}[\tilde{z}^{1},\dots,\tilde{z}^{\ell}$], then for each $i=1,\dots,\ell$ we have
$z^{i}-\tilde{z}^{i}=\lambda(s^{i}-\tilde{s}^{i}).$ (2.42)
Here we write $\tilde{z}^{i}=y^{i}+\lambda\tilde{s}^{i}$. Since $z^{i}-\tilde{z}^{i}\in\mathscr{P}=\mathbb{C}[z^{1},\dots,z^{\ell}]$ and $(s^{i}-\tilde{s}^{i})\in\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, if $s^{i}-\tilde{s}^{i}$ do not vanish, then (2.42) gives a nontrivial algebraic relation of $z^{1},\dots,z^{\ell}$ and $\lambda$, which contradicts with their algebraic independence. Therefore, $z^{i}=\tilde{z}^{i}$ for each $i=1,\dots,\ell$. ∎
The above theorem shows that the map
$\{\text{proper~generators}\}\to\{\text{proper~factors}\},\quad\{z^{1},\dots,z^{\ell}\}\mapsto\mathbb{C}[z^{1},\dots,z^{\ell}].$ (2.43)
is a bijection. As a result, we can characterize a proper factor by its proper generators.
For any given set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we still call the transformation which maps $(x^{1},\dots,x^{\ell})$ to $(z^{1},\dots,z^{\ell})$ the $\varphi_{\lambda}$-transformation, and the corresponding $\varphi_{0}$-transformation the null-transformation. Note that $\{z^{1},\dots,z^{\ell}\}$ and $\{y^{1},\dots,y^{\ell}\}$ may lead to distinct $\varphi_{\lambda}$-transformations, but they always share the same null-transformation. In particular, for any element $r$ of the set $S$ defined by (2.29) we have
@@ 3. The construction of generalized Frobenius manifold structures
For a given irreducible reduced root system $R$ in an $\ell$-dimensional Euclidean space $V$, let us fix a weight $\omega$ of the form (2.1), a proper factor $\mathscr{P}$ of $\mathscr{A}^{W}$ with the corresponding set $\{z^{1},\dots,z^{\ell}\}$ of proper generators. We are to impose a certain additional condition on these proper generators, so that we construct a generalized Frobenius manifold structure on a dense open subset of the spectrum of $\mathscr{P}$.
@@ 3.1. The orbit space and pencil generators
Consider the manifold $\mathcal{M}=\mathbb{C}^{\ell}$ on which $z^{1},\dots,z^{\ell}$ are global coordinates. Taking $x^{1},\dots,x^{\ell}$ as coordinates on the complexified Euclidean space $V\otimes\mathbb{C}$, we have a collection of holomorphic maps
@@
Here the $\varphi_{\lambda}$-transformation is defined in the above section. We call $\mathcal{M}$ the orbit space of the affine Weyl group $W_{a}(R)$ w.r.t. the proper generators $z^{1},\dots,z^{\ell}$. In this section, we will use this collection of maps $\{\varphi_{\lambda}\}_{\lambda\in\mathbb{C}}$ to introduce a structure of generalized Frobenius manifold on a dense open subset of $\mathcal{M}$.
The inner product (,)($\cdot\,$$\cdot$) of the Euclidean space $V$ yields a contravariant metric
$a(\mathrm{d}x^{i},\mathrm{d}x^{j})=a^{ij}\quad\mbox{with}\ (a^{ij})=\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)^{-1},\quad i,j=1,\dots,\ell$ (3.2)
on $V\otimes\mathbb{C}$. By a contravariant metric, we mean a non-degenerate, symmetric bilinear form on the holomorphic cotangent bundle $T^{*}(V\otimes\mathbb{C})$. We will also call $a$ a metric on $V\otimes\mathbb{C}$.
@@
Note that we introduce the factor $1/{4\pi^{2}}$ to simplify the expressions of $g_{\lambda}^{ij}$’s. We denote
$\Gamma^{ij}_{g_{\lambda},k}=-g_{\lambda}^{ir}\Gamma_{g_{\lambda},rk}^{j},\quad i,j,k=1,\dots,\ell,$ (3.6)
and call them the Christoffel symbols of the Levi-Civita connection of the contravariant metric $g_{\lambda}$, or simply the Christoffel symbols of $g_{\lambda}$.
@@
For all indices $i,j,k$, we have $g^{ij}_{\lambda}\in\mathscr{A}^{W}$ and $\Gamma^{ij}_{g_{\lambda},k}\in\mathscr{A}^{W}$.
@@
We first consider the entries of the contravariant metric $g_{\lambda}$. By definition, we have $g^{ij}_{\lambda}\in\mathscr{A}$, hence it suffices to show that $g^{ij}_{\lambda}$ are invariant w.r.t. the action of $W_{a}(R)$. Since the invariance of $g^{ij}_{\lambda}$ w.r.t. translations is obvious, we consider actions by elements of the Weyl group $W(R)$. Write $z^{j}$ as $z^{j}=z^{j}(x^{1},\dots,x^{\ell};c)$. For any given $\sigma\in W(R)$, since $z^{j}(x^{1},\dots,x^{\ell};c)=z^{j}(\sigma(x^{1},\dots,x^{\ell};c))$, we have
@@
$\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}=a^{pq},$
hence we arrive at $g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=g^{ij}_{\lambda}(\sigma(x^{1},\dots,x^{\ell};c))$.
Now we consider the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$ of the contravariant metric $g_{\lambda}$. Since $g_{\lambda}$ is invariant w.r.t. the action of $W_{a}(R)$, it suffices to show that $\Gamma^{ij}_{g_{\lambda},k}\in\mathscr{A}$. From Corollary 2.14 and Proposition 2.6, it follows that
$\det\left(\frac{\partial z^{i}}{\partial x^{j}}\right)=\mathbf{J}(\hat{\mathbf{x}};\lambda)=C\cdot\mathrm{e}^{\pi i\sum_{\alpha\in R^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in R^{+}}\left(1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}\right),$
where $\mathbf{x}=c\omega+\hat{\mathbf{x}}$. For the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$, the formula
$\Gamma^{ij}_{g_{\lambda},k}=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{p}}\frac{\partial^{2}z^{j}}{\partial x^{q}\partial x^{r}}a^{pq}\frac{\partial x^{r}}{\partial z^{k}}$ (3.8)
holds, hence we have
$\Gamma^{ij}_{g_{\lambda},k}=\frac{P^{ij}_{k}(\hat{\mathbf{x}};\lambda)}{\mathbf{J}(\hat{\mathbf{x}};\lambda)},$
where $P_{k}^{ij}\in\mathscr{A}$. Note that $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ is anti-invariant and $\Gamma_{\lambda,k}^{ij}$ are invariant w.r.t. the action of the Weyl group $W(R)$, so the Fourier polynomials $P_{k}^{ij}(\hat{\mathbf{x}};\lambda)$ must be anti-invariant w.r.t. the action of $W(R)$. Note that for each $\beta\in R^{+}$, if $(\beta,\mathbf{x})\in\mathbb{Z}$,
$\sigma_{\beta}(\mathbf{x})=\mathbf{x}-(\beta,\mathbf{x})\beta^{\vee}=\mathbf{x}+n_{1}\alpha_{1}^{\vee}+\cdots+n_{\ell}\alpha_{\ell}^{\vee},\quad n_{1},\dots,n_{\ell}\in\mathbb{Z}.$ (3.9)
Here $\beta^{\vee}\in R^{\vee}$ is a $\mathbb{Z}$-combination of $\alpha_{1}^{\vee},\dots,\alpha_{\ell}^{\vee}$. It follows from the anti-invariance of $P_{k}^{ij}$ and (3.9) that, as long as $(\beta,\mathbf{x})\in\mathbb{Z}$,
$P_{k}^{ij}(\mathbf{x})=-P_{k}^{ij}(\sigma_{\beta}(\mathbf{x}))=-P_{k}^{ij}(\mathbf{x}),$
hence $P_{k}^{ij}$ vanishes along $(\beta,\mathbf{x})\in\mathbb{Z}$, all zeros of $1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}$. Since $1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}$ has no square factor in $\mathscr{A}$, it follows from Nullstellensatz’s Theorem that $P_{k}^{ij}$ is divisible by $1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}$. Therefore, it is also divisible by $\mathbf{J}(\hat{\mathbf{x}};\lambda)$, and we have $\Gamma^{ij}_{g_{\lambda},k}\in\mathscr{A}$. The proposition is proved. ∎
Remark 3.1.
This proposition implies that the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$ can be defined globally on $\mathcal{M}$, although $g_{\lambda}$ may degenerate at some points.
The above proposition shows, in particular, that $g^{ij}_{\lambda}$ and $\Gamma_{g_{\lambda},k}^{ij}$ are polynomials of $\lambda$. In order to construct generalized Frobenius manifold structures on a dense open subset of the orbit space $\mathcal{M}$, we are to impose an additional condition on the proper generators of $\mathscr{A}^{W}$, so that $g_{\lambda}$ yields a flat pencil of metrics [10] on $\mathcal{M}$.